During a football game, the parents of the football players sell pretzels and popcorn to raise money for the new uniforms. They charge $2.50 for a bag of popcorn $2 for a pretzel. The parents collect $336 in sales during the game. They sell twice as many bags of popcorn as pretzels. How many bags of popcorn do they sell? How many pretzels do they sell?
Answer by checkley79(3341)

X=2Y
2.50X+2Y=336
2.50*2Y+2Y=336
5Y+2Y=336
7Y=336
Y=336/7
Y=48 BAGS OF PRETZELS WERE SOLD.
X=2*48=96 BAGS OF POPCORN WERE SOLD.
PROOF:
2.50*96+2*48=336
240+96=336
336=336